OpenStudy (abmon98):

Prove the identity sinx/1+cosx=1-cosx/sinx

OpenStudy (luigi0210):

Multiply by the conjugate

OpenStudy (abmon98):

\[Sinx \div1+cosx, Sinx/1+sinx/cosx, sinx*cosx+sinx/cosx\]

OpenStudy (alekos):

you haven't multiplied the numerator and denominator by the conjugate 1-cosx

OpenStudy (anonymous):

\[\frac{ \sin x }{ 1+\cos x } \times \frac{ 1-\cos x }{ 1-\cos x }\] \[= \frac{ \sin x(1-\cos x) }{(1^{2}-\cos ^{2}x) }\] \[=\frac{ \sin x(1-\cos x) }{ \sin ^{2}x }=\frac{ (1-\cos x) }{ \sin x }\]